OpenStudy (anonymous):

Which of the following is the slope between the points (-1, -5) and (7, -3)?

OpenStudy (anonymous):

(-1,-5) (7,-3) Change of x: 8 Change of y: 2 slope = Change of y/Change of x = 2/8 = 1/4 Where are your 'followings'?

OpenStudy (anonymous):

1/4

OpenStudy (anonymous):

basically ur standarad equation is y=ax+b first u have to find a which is equal to (y2-y1)/(x2-x1) and then just solve for ur b by plugging in a and either point :) tahsts what we've been doing for ur other questions to :)